OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the given absolute value inequality? |x + 5| > 18 I got this am I right or wrong? x + 5 > or x + 5 < -18 Which is the solution set to the given inequality? |x + 3| < 12 I got -15 u 9 am i right or wrong? HELP PLZ!

OpenStudy (anonymous):

\[\left[\begin{matrix}x>-2 & \\ x+y<4& \end{matrix}\right]\] NEED HELP ON THIS ONE TO

OpenStudy (aum):

|x + 5| > 18 implies x + 5 > 18 OR x + 5 < -18

OpenStudy (aum):

x + 5 > 18 x > 13 OR x + 5 < -18 x < -23 \(x = (-\infty, -23) \cup (13, \infty)\)

OpenStudy (aum):

|x + 3| < 12 implies x + 3 < 12 AND x + 3 > -12 x + 3 < 12 x < 9 AND x + 3 > -12 x > -15 Solution: -15 < x < 9 In interval notation: (-15, 9).

OpenStudy (anonymous):

would it be written like this? @aum

OpenStudy (anonymous):

wouldnt it be [-15,9]

OpenStudy (aum):

No, the end points -15 and 9 are not included in the solution -15 < x < 9 and so it should be (-15,9). If the solution had been \(-15 \le x \le 9\) then the solution would be [-15. 9] where both end points are included in the solution. So for this problem the solution is: (-15, 9).

OpenStudy (anonymous):

oh I see thanks!!!! can you help me on this last one plz \[\left[\begin{matrix}y >-2 & \\ x+y<4& \end{matrix}\right]\]